OpenStudy (anonymous):

some Projectiles are thrown with 0

OpenStudy (anonymous):

state your answer in terms of \[a=\frac{v_{0}^2}{4g}\]

OpenStudy (anonymous):

is the answer \[s = -\nu _{o}^{2}\sin ^{2}\theta / 2g\] ?

OpenStudy (anonymous):

thats right \[y=\frac{v_{0}^2 \sin^2 \theta}{2g}\] but can u change it to x,y (cartesian coordinates)

OpenStudy (anonymous):

@Vaidehi09 \[\sin \theta = \frac{y}{\sqrt{x^2+y^2}}\]

OpenStudy (anonymous):

oh. yes u can do that. that'll be just playing with the terms.